# Diffusion Policy Policy Optimization

### v1

@@ Appendix A Extended Related Work
This section presents an extended discussion on related methods that train/improve diffusion-based policies with RL-related methods. The baselines to which we compare in Section 5.1 are discussed below as well, and are highlighted in their corresponding colors. We also refer the readers to Zhu et al. [89] for an extensive survey on diffusion models for RL.
@@
Most previous work have focused on the offline setting with a static dataset. One line of work focuses on state trajectory planning and guiding the denoising sampling process such that the sampled actions satisfy some desired objectives. Janner et al. [35] applies classifier guidance that generates trajectories with higher predicted reward. Ajay et al. [2] introduces classifier-free guidance that avoids learning the value of noisy states. There is another line of work that uses diffusion model as action policy (instead of state planner) and generally applies Q-learning. DQL [84] introduces Diffusion Q-Learning that learns a state-action critic for the final denoised actions and backpropagates the gradient from the critic through the entire Diffusion Policy (actor), akin to the usual Q-learning. IDQL [27], or Implicit Diffusion Q-learning, proposes learning the critic to select the actions at inference time for either training or evaluation, while fitting the actor to all sampled actions. Kang et al. [37] instead proposes using the critic to re-weight the sampled actions for updating the actor itself, similar to weighted regression baselines DAWR and DRWR introduced in our work. Goo and Niekum [26] similarly extracts the policy in the spirit of AWR [53]. Chen et al. [12] trains the critic using value iteration instead based on samples from the actor.
@@
We note that methods like DQL and IDQL can also be applied in the online setting. There is also a small amount of work that focuses entirely on the online setting. DIPO [86] differs from DQL and related work that it uses the critic to update the sampled actions (“action gradient”), instead of the actor — the actor is then fitted with updated actions from the replay buffer. QSM, or Q-Score Matching [59], suggests that optimizing the likelihood of the entire chain of denoised actions can be inefficient (contrary to our findings in the fine-tuning setting), and instead proposes learning the optimal policy by iteratively aligning the gradient of the actor (i.e., score) with the action gradient of the critic. Rigter et al. [64] proposes learning a diffusion dynamic model to generate synthetic trajectories for training a non-diffusion RL policy online.
Appendix B Additional experimental results
B.1 Ablation studies on design decisions in DPPO
@@
Fig. A1 shows the fine-tuning results in Hopper-v2 and HalfCheetah-v2 from Gym, and Can and Square from Robomimic. On the simpler Hopper-v2, we observe that the two baselines, both estimating the value of some action, achieves higher reward during fine-tuning than DPPO’s choice. However, in the more challenging tasks, the environment-state-only advantage used in DPPO consistently leads to the most improved performance. We believe estimating the accurate value of applying a continuous and high-dimensional action can be challenging, and this is exacerbated by the high stochasticity of diffusion-based policies and the action chunk size. The results here corroborate the findings in Section 5.1 that off-policy Q-learning methods can perform well in Hopper-v2 and Walker2D-v2, but exhibit training instability in manipulation tasks from Robomimic.
@@
Figure A3: Choice of minimum diffusion noise. Results are averaged over three seeds.
@@
B.2 Effect of expert data
@@
In Fig. A6 we compare DPPO (10 denoising steps) and Gaussian trained from scratch (no pre-training on expert data) in the three OpenAI Gym tasks. As using larger action chunk sizes $T_{a}$ leads to poor from-scratch training shown in Fig. A5, we focus on single-action chunks $T_{a}=1$ as is typical in RL benchmarking. Though we find Gaussian trains faster than DPPO (expected since DPPO solves an MDP with longer effective horizon), DPPO still attains reasonable final performance. However, due to the multi-step (10) denoising sampling, DPPO takes about 6×6$\times$6 × wall-clock time compared to Gaussian. We hope that future work will explore how to design the training curriculum of denoising steps for the best balance of training performance and wall-clock efficiency.
@@
B.3 Comparing to other policy parameterizations in Avoid
@@
B.4 Comparing to other policy parameterizations in the easier tasks from Robomimic
@@
B.5 Comparing to policy gradient using exact likelihood of Diffusion Policy
@@
Fig. A9 shows the comparison between DPPO and diffusion policy gradient using exact likelihood estimate. Exact policy gradient improves the base policy in Hopper-v2 but does not outperform DPPO. It also requires more runtime and GPU memory as it backpropagates through the ODE. In the more challenging Can its success rate drops to zero. Moreover, policy gradient with exact likelihood does not offer the flexibility of fine-tuning fewer-than-$K$ denoising steps or discounting the early denoising steps that DPPO offers, which have shown in Section B.1 to often improve fine-tuning efficiency.
@@
Appendix C Reporting of Wall-Clock Times
@@
Table A1 and Table A2 shows the the wall-clock time used in each OpenAI Gym task and Robomimic task. In Gym tasks, on average DPPO trains 74%, 41%, 37%, and 12% faster than QSM, DAWR, DIPO, and DQL, respectively, which all require a significant amount of gradient updates per sample to train stably. DRWR and IDQL trains 33% and 7% faster than DPPO, respectively. Robomimic tasks are more expensive to simulate, especially with Transport task, and thus the wall-clock difference is smaller among the different methods. On average DPPO trains 29%, 3%, 7%, and 3% faster than QSM, DAWR, DIPO, and DQL, respectively. DRWR trains 3% faster than DPPO, and IDQL uses comparable time as DPPO.
@@
Comparison of other policy parameterizations and architecture Section 5.2 and Section 5.3.
@@
Appendix D Additional Experimental Details
D.1 Details of policy architectures used in all experiments
MLP.
Transformer.
UNet.
ViT.
Task Obs dim - State Obs dim - Pixel Act dim $T$ Sparse reward ?
@@ D.1 Details of policy architectures used in all experiments
D.2 Additional details of Gym tasks and training in Section 5.1

### v3

@@ A.2 Diffusion-based RL methods
This section discusses related methods that directly train or improve diffusion-based policies with RL methods. The baselines to which we compare in Section 5.1 are discussed below as well, and are highlighted in their corresponding colors. We also refer the readers to Zhu et al. [108] for an extensive survey on diffusion models for RL.
@@
Most previous works have focused on the offline setting with a static dataset. One line of work focuses on state trajectory planning and guiding the denoising sampling process such that the sampled actions satisfy some desired objectives. Janner et al. [45] apply classifier guidance that generates trajectories with higher predicted rewards. Ajay et al. [2] introduce classifier-free guidance that avoids learning the value of noisy states. There is another line of work that uses diffusion models as an action policy (instead of state planner) and generally applies Q-learning. DQL [100] introduces Diffusion Q-Learning that learns a state-action critic for the final denoised actions and backpropagates the gradient from the critic through the entire Diffusion Policy (actor) denoising chain, akin to the usual Q-learning. IDQL [35], or Implicit Diffusion Q-learning, proposes learning the critic to select the actions at inference time for either training or evaluation while fitting the actor to all sampled actions. Kang et al. [47] instead propose using the critic to re-weight the sampled actions for updating the actor itself, similar to weighted regression baselines DAWR and DRWR introduced in our work. Goo and Niekum [31] similarly extract the policy in the spirit of AWR [67]. Chen et al. [16] train the critic using value iteration instead based on samples from the actor. Finally, Jackson et al. [44] explore using diffusion guidance to move offline data towards the target trajectory distribution.
@@
We note that methods like DQL and IDQL can also be applied in the online setting. A small amount of work also focuses entirely on the online setting. DIPO [103] differs from DQL and related work in that it uses the critic to update the sampled actions (“action gradient”) instead of the actor — the actor is then fitted with updated actions from the replay buffer. QSM, or Q-Score Matching [74], suggests that optimizing the likelihood of the entire chain of denoised actions can be inefficient (contrary to our findings in the fine-tuning setting) and instead proposes learning the optimal policy by iteratively aligning the gradient of the actor (i.e., score) with the action gradient of the critic. Rigter et al. [79] proposes learning a diffusion dynamic model to generate synthetic trajectories for online training of a non-diffusion RL policy.
We note that almost all prior work in diffusion-based RL (offline or online) have relied on approximating the state-action Q function and using it to update the diffusion actor in some form — policy gradient update has been deemed challenging due to the multi-step denoising process [74, 103]. Inaccurate Q values may lead to biased updates to the actor, which can lead to training collapse as it starts with decent pre-training performance but quickly drops to zero success rate as seen in Fig. 5, also failing to recover since then due to the sparse-reward setup. While Q-learning methods generally achieve better sample efficiency when they can solve the task of interest, our focus has been largely on challenging long-horizon robot manipulation tasks where the training stability is much desired.
Distinction from the policy gradient formulation in Psenka et al. [74].
There has been a different formulation introduced in Psenka et al. [74] Sec. 3 that derives the policy gradient update for diffusion policy. The derivation is based on converting the gradient of the log likelihood of the final denoised action to the sum over log likelihood of individual denoising actions. This formulation, unlike DPPO, does not treat the multi-step denoising process as a MDP. In the policy gradient update, Psenka et al. [74] sum over denoising steps and then takes expectation over environment steps, while DPPO’s update (4.4) takes expectation over both denoising and environment steps, potentially leading to better sample efficiency. Moreover, Psenka et al. [74] do not propose applying PPO updates or other modifications to diffusion, and finds such vanilla form of policy gradient update to be ineffective. We formulate DPPO independent of their work and find DPPO highly effective in fine-tuning settings while also being competitive in training from scratch (Section C.3).
Appendix B Additional details of DPPO implementation
Pseudocode.
The pseudocode for DPPO is presented in Algorithm 1. DPPO takes as input a diffusion policy $\pi_{\theta}$ trained using behavior cloning loss $\mathcal{L}_{\text{BC}}$. The policy is then fine-tuned using a PPO-style loss [85] with careful treatment of the denoising process (Section 4).
Pre-training.
The diffusion policy $\pi_{\theta}$ is pre-trained using a behavior cloning loss [38]:
$\mathcal{L}_{\text{BC}}(\theta)=\mathbb{E}^{(s_{t},a_{t}^{0})\sim\mathcal{D}_{\text{off}}}\big{[}\|\varepsilon_{t}-\varepsilon_{\theta}(a_{t}^{0},s_{t},k)\|^{2}\big{]},$ (A1)
where $\mathcal{D}_{\text{off}}$ is the offline dataset and $\varepsilon_{\theta}$ is the policy network predicting the sampled noise added to $a_{t}^{0}$ based on the noisy action. We use the cosine noise schedule from Nichol and Dhariwal [63].
Algorithm 1 DPPO
1:Pre-train diffusion policy $\pi_{\theta}$ with offline dataset $\mathcal{D}_{\text{off}}$ using BC loss $\mathcal{L}_{\text{BC}}(\theta)$ Eq. (A1).
2:Initialize value function $V_{\phi}$.
3:for iteration = 1, 2, … do
4:     Initialize rollout buffer $\mathcal{D}_{\text{itr}}$.
5:     $\pi_{\theta_{\text{old}}}=\pi_{\theta}$.
6:     for environment = 1, 2, …, $N$ in parallel do
7:         Initialize state $\bar{s}_{\bar{t}(0,K)}=(s_{0},a_{0}^{K})$ in $\mathcal{M}_{\textsc{DP}}$.
8:         for environment step t = 1, …, $T$, denoising step $k=K-1,\dots,0$ do
9:              Sample the next denoised action $\bar{a}_{\bar{t}(t,k)}=a_{t}^{k}\sim\pi_{\theta_{\text{old}}}$.
10:              if $k=0$ then
11:                  Run $a_{t}^{0}$ in the environment and observe $\bar{R}_{\bar{t}(t,0)}$ and $\bar{s}_{\bar{t}(t+1,K)}$.
12:              else
13:                  Set $\bar{R}_{\bar{t}(t,k)}=0$ and $\bar{s}_{\bar{t}(t,k-1)}=(s_{t},a_{t}^{k})$.
14:              Add $(k,\bar{s}_{\bar{t}(t,k)},\bar{a}_{\bar{t}(t,k)},\bar{R}_{\bar{t}(t,k)})$ to $\mathcal{D}_{\text{itr}}$.
15:     Compute advantage estimates $A^{\pi_{\theta_{\text{old}}}}(s_{\bar{t}(t,k=0)},a_{\bar{t}(t,k=0)})$ for $\mathcal{D}_{\text{itr}}$ using GAE Eq. (A2).
[…]
17:         for minibatch = 1, 2, …, $B$ do
18:              Sample $(k,\bar{s}_{\bar{t}(t,k)},\bar{a}_{\bar{t}(t,k)},\bar{R}_{\bar{t}(t,k)})$ and $A^{\pi_{\theta_{\text{old}}}}(s_{\bar{t}(t,k)},a_{\bar{t}(t,k)})$ from $\mathcal{D}_{\text{itr}}$.
19:              Compute denoising-discounted advantage $\hat{A}_{\bar{t}(t,k)}=\gamma_{\textsc{denoise}}^{k}A^{\pi_{\theta_{\text{old}}}}(s_{\bar{t}(t,0)},a_{\bar{t}(t,0)})$.
20:              Optimize $\pi_{\theta}$ using policy gradient loss $\mathcal{L}_{\theta}$ Eq. (A3).
21:              Optimize $V_{\phi}$ using value loss $\mathcal{L}_{\phi}$ Eq. (A4).
22:return converged policy $\pi_{\theta}$.
Environment-step advantage estimation.
We use Generalized Advantage Estimation (GAE) [84] with parameter $\lambda$ for advantage estimation in Algorithm 1. GAE-$\lambda$ approximates the advantage function using the series
$\hat{A}_{\bar{t}(t,k=0)}^{\lambda}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\bar{\delta}_{\bar{t}(t+l,k=0)},\quad\text{where }\bar{\delta}_{\bar{t}(t,k)}=\bar{R}_{\bar{t}(t,k)}+\gamma_{\text{ENV}}V_{\phi}(\bar{s}_{\bar{t}(t+1,k)})-V_{\phi}(\bar{s}_{\bar{t}(t,k)}).$ (A2)
Notably, GAE-$\lambda$ interpolates between a one-step temporal difference ($\hat{A}_{\bar{t}(t,k)}^{\lambda=0}=\bar{R}_{\bar{t}(t,k)}+\gamma_{\text{ENV}}V_{\phi}(\bar{s}_{\bar{t}(t+1,k)})-V_{\phi}(\bar{s}_{\bar{t}(t,k)})$) and the Monte Carlo return of the episode relative to the baseline ($\hat{A}_{\bar{t}(t,k)}^{\lambda=1}=\sum_{l=0}^{T-t}\gamma_{\text{ENV}}^{l}\bar{R}_{\bar{t}(t+l,k)}-V_{\phi}(\bar{s}_{\bar{t}(t,k)})$). We refer the reader to Table A8 for additional details on GAE parameter selection and Section C.2 for ablations on the choice of advantage estimator.
Note that in Eq. (A2) we are only concerned with $k=0$, i.e., the final denoising step, calculating the advantage for $k=0$ (i.e., environment steps) but not for intermediate denoising steps. We only need to apply denoising discounting to the calculated advantages so they can be applied to each denoising step $k$.
Fine-tuning.
During RL fine-tuning, we update the policy $\pi_{\theta}$ using the clipped objective:
$\mathcal{L}_{\theta}=\mathbb{E}^{\mathcal{D}_{\text{itr}}}\min\Big{(}\hat{A}^{\bar{\pi}_{\theta_{\text{old}}}}(\bar{s}_{\bar{t}},\bar{a}_{\bar{t}})\frac{\bar{\pi}_{\theta}(\bar{s}_{\bar{t}},\bar{a}_{\bar{t}})}{\bar{\pi}_{\theta_{\text{old}}}(\bar{s}_{\bar{t}},\bar{a}_{\bar{t}})},\ \hat{A}^{\pi_{\theta_{\text{old}}}}(\bar{s}_{\bar{t}},\bar{a}_{\bar{t}})\operatorname{clip}\big{(}\frac{\bar{\pi}_{\theta}(\bar{s}_{\bar{t}},\bar{a}_{\bar{t}})}{\bar{\pi}_{\theta_{\text{old}}}(\bar{s}_{\bar{t}},\bar{a}_{\bar{t}})},1-\varepsilon,1+\varepsilon\big{)}\Big{)}.$ (A3)
If we choose to fine-tune only the last $K^{\prime}$ denoising steps, then we sample only those from $\mathcal{D}_{\text{itr}}$.
Finally, we train the value function to predict the future discounted sum of rewards (i.e., discounted returns):
$\mathcal{L}_{\phi}=\mathbb{E}^{\mathcal{D}_{\text{itr}}}\big{[}\|\sum_{l=0}^{T-t}\gamma_{\textsc{env}}^{l}\bar{R}_{\bar{t}(t+l,k)}-V_{\phi}(s_{t})\|^{2}\big{]}.$ (A4)
Similar to all baselines in Section E.3, we denote $N_{\theta}$ and $N_{\phi}$ the replay ratio for the actor (diffusion policy) and the value critic in DPPO; in practice we always set $N_{\theta}=N_{\phi}$. Similar to usual PPO implementations [41], the batch updates in an iteration terminate when the KL divergence between $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ reaches 1, although in practice we find this never happens.
Large batch size.
Since the gradient update in DPPO involves expectation over both environment steps and denoising steps, we use a larger batch size compared to, e.g., PPO training with Gaussian policy parameterization. Roughly we use the batch size from Gaussian training times the number of the fine-tuned denoising steps; in some cases like Robomimic we also observe that a much smaller batch size (close to that of Gaussian training) can be used and significantly improves sample efficiency.
Gradient clipping ratio.
[…]
Appendix C Additional experimental results
C.1 Comparing to demo-augmented RL baselines using diffusion policy instead
In Section 5.2 we compare DPPO with other demo-augmented RL methods, namely, RLPD, Cal-QL, and IBRLDPPO uses diffusion policy while the baselines use Gaussian policy. Here we experiment with using diffusion policies for the baselines and the results are shown in Fig. A1. We use either action chunk size $T_{a}=1$ or $T_{a}=4$. We see similar results as in Fig. 6 using Gaussian policies that RLPD and Cal-QL fails to solve the task at all. We believe that the worse performance of Cal-QL is due to the offline RL objective (based on learning the state-action Q function) making learning precise continuous actions needed in Robomimic tasks very difficult, regardless of the policy parameterization, which corroborates our original finding in Section 5.1 when comparing DPPO to Q-learning-based diffusion RL methods. Compared to RLPD that trains with the SAC objective and expert data in the replay buffer, IBRL, using BC pre-training, is able to learn a base policy more effectively and uses it for online data collection. DPPO benefits from directly fine-tuning the pre-trained policy (instead of training a new one using experiences from the pre-trained policy), and achieves similar or better sample efficiency before 1e6 steps compared to IBRL, and converges to $\sim$100success rates unlike IBRL saturates at lower levels (not shown).
Figure A1: Using diffusion policy for other demo-augmented RL methods. Results are averaged over three seeds.
C.2 Ablation studies on design decisions in DPPO
@@
Fig. A2 shows the fine-tuning results in Hopper-v2 and HalfCheetah-v2 from Gym, and Can and Square from Robomimic. On the simpler Hopper-v2, we observe that the two baselines, both estimating the value of some action, achieves higher reward during fine-tuning than DPPO’s choice. However, in the more challenging tasks, the environment-state-only advantage used in DPPO consistently leads to the most improved performance. We believe estimating the accurate value of applying a continuous and high-dimensional action can be challenging, and this is exacerbated by the high stochasticity of diffusion-based policies and the action chunk size. The results here corroborate the findings in Section 5.1 that off-policy Q-learning methods can perform well in Hopper-v2 and Walker2D-v2, but often exhibit training instability in manipulation tasks from Robomimic.
@@
Figure A4: Choice of minimum diffusion noise. Results are averaged over three seeds. Note in Left, with higher minimum noise level, the sampled trajectories exhibit wider coverage at the two modes but still maintain the overall structure.
@@
C.3 Effect of expert data
@@
In Fig. A7 we compare DPPO (10 denoising steps) and Gaussian trained from scratch (no pre-training on expert data) in the three OpenAI Gym tasks. As using larger action chunk sizes $T_{a}$ leads to poor from-scratch training shown in Fig. A6, we focus on single-action chunks $T_{a}=1$ (and $T_{p}=1$) as is typical in RL benchmarking. Though we find Gaussian trains faster than DPPO (expected since DPPO solves an MDP with longer effective horizon), DPPO still attains reasonable final performance. However, due to the multi-step (10) denoising sampling, DPPO takes about 6×6$\times$6 × wall-clock time compared to Gaussian. We hope that future work will explore how to design the training curriculum of denoising steps for the best balance of training performance and wall-clock efficiency.
@@
C.4 Comparing to other policy parameterizations in Avoid
@@
C.5 Comparing to other policy parameterizations in the easier tasks from Robomimic
@@
C.6 Comparing to policy gradient using exact likelihood of Diffusion Policy
@@
Fig. A10 shows the comparison between DPPO and diffusion policy gradient using exact likelihood estimate. Exact policy gradient improves the base policy in Hopper-v2 but does not outperform DPPO. It also requires more runtime and GPU memory as it backpropagates through the ODE. In the more challenging Can its success rate drops to zero. Moreover, policy gradient with exact likelihood does not offer the flexibility of fine-tuning fewer-than-$K$ denoising steps or discounting the early denoising steps that DPPO offers, which have shown in Section C.2 to often improve fine-tuning efficiency.
@@
C.7 Ablating Structured Exploration in DPPO
Here we provide additional evidence on how structured exploration of DPPO (Section 6) aids RL fine-tuning. While Fig. A8 compares DPPO with Gaussian and GMM policies and shows DPPO trajectories achieve wide coverage and stay near the expert data manifold, in Fig. A11 we ablate such structured exploration within DPPO. We use DDIM [87] such that actions can be sampled deterministically — this allows us to sample trajectories without adding any noise to intermediate denoising steps but only to the final denoised action ($k=0$), and compare that to DPPO with noise added to all denoising steps. In both cases, we consider the minimum noise level $\sigma_{\text{min}}^{\text{exp}}$ of 0.05 and 0.1. We see in Fig. A11 that with higher noise level, DPPO trajectories cover the expert data modes well without exploring aggressively into new modes, while in the case of only adding noise to the final step, the trajectories become less structured especially in M3.
Then we run both exploration schemes in Can and Square from Robomimic, and Fig. A3 right shows the original DPPO setup achieves faster convergence than when noise is only added to the final step. This result, on top of results from Section 5.3 showing DPPO achiving better sample efficiency than Gaussian and GMM policies, showcases the benefit of structured exploration in fine-tuning.
Figure A11: Ablating Structured Exploration in DPPO. (Left) Sampled trajectories with noise added to all denoising steps vs. only to the last step $k=0$ in Avoid. (Right) Results are averaged over three seeds in Can and Square.
Appendix D Reporting of Wall-Clock Times
@@
Table A1 and Table A2 shows the the wall-clock time used in each OpenAI Gym task and Robomimic task. In Gym tasks, on average DPPO trains 41%, 37%, and 12% faster than DAWR, DIPO, and DQL, respectively, which all require a significant amount of gradient updates per sample to train stably. QSM, DRWR, and IDQL trains 43%, 33%, and 7% faster than DPPO, respectively. Robomimic tasks are more expensive to simulate, especially with Transport task, and thus the wall-clock difference is smaller among the different methods. All methods use comparable time except for DIPO that uses slightly more on average.
@@
Comparing to other policy parameterizations and architecture Section 5.3 and Section 5.4.
@@
Appendix E Additional Experimental Details
Task / Dataset Obs dim - State Obs dim - Pixel Act dim $T$ Sparse reward ?
@@ Appendix E Additional Experimental Details
Franka-Kitchen Kitchen-Complete-v0 60 - 9 280 Yes
Kitchen-Partial-v0 60 - 9 280 Yes
Kitchen-Mixed-v0 60 - 9 280 Yes
@@
E.1 Details of policy architectures used in all experiments
MLP.
Transformer.
UNet.
ViT.
E.2 Additional details of Gym tasks and training in Section 5.1